# EEPPR: Event-based Estimation of Periodic Phenomena Rate using Correlation in 3D

### v2

@@ Abstract
We present a novel method for measuring the period of phenomena like rotation, flicker and vibration, by an event camera, a device asynchronously reporting brightness changes at independently operating pixels with high temporal resolution. The approach assumes that for a periodic phenomenon, a highly similar set of events is generated within a spatio-temporal window at a time difference corresponding to its period. The sets of similar events are detected by a correlation in the spatio-temporal event stream space. The proposed method, EEPPR, is evaluated on a dataset of $12$ sequences of periodic phenomena, i.e. flashing light and vibration, and periodic motion, e.g., rotation, ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM). EEPPR significantly outperforms published methods on this dataset, achieving the mean relative error of $0.1\%$. The dataset and codes are available on GitHub.
@@ 1 Introduction
Traditional methods for measuring periodic phenomena often involve contact-based devices such as mechanical tachometers and rotary encoders. These approaches, while established, have inherent limitations: (i) physical contact with the target object can influence its movement and introduce measurement inaccuracies, (ii) contact methods may not be feasible in scenarios with delicate objects, confined spaces, or situations where interfering with the target significantly affects its movement. Laser tachometers offer a less invasive and highly accurate [^1] See https://meters.uni-trend.com/product/ut370-series/#Specifications. alternative but require reflective material (e.g., a sticker) on the target and precise aiming of the laser, as missing the reflective material pass-through results in an inaccurate measurement. Although contactless event-based methods exist, they are either handcrafted for a specific type of periodic phenomena [1] or require user supervision [13, 14].
Figure 1: EEPPR method: (i) data captured from an event camera is aggregated into a 3D array, (ii) the array is split into same-sized areas, and in each area a template depth is automatically selected, (iii) a 3D correlation of the template with the event stream is computed, (iv) a frequency is calculated from the median of time deltas measured between correlation peaks for each window, (v) the output frequency is computed as a median of measurements from all windows.
In this paper, we propose a non-contact method for measuring the properties of any periodic phenomena using an event camera. The input of our method is the event stream, and the output is a single scalar – the rate of the periodic phenomena. The method operates remotely, without requiring any modification of the observed object (e.g. a marker). The approach assumes that similar sets of pixel activations of an event camera (events) will occur at regular time intervals in any sequence capturing a periodic phenomenon. These intervals correspond to the period.
The method was evaluated on a new dataset of twelve sequences of four types of periodic phenomena: (i) flickering light, (ii) object vibration, (iii) rotation, and (iv) repetitive translational movement. A mean relative error of $0.1\%$ was achieved, which is 140$\approx$ 140$\times$ 140 × lower than the mean relative error of the best available published method (see Tab. 3). The point estimate for the period lengths is obtained by calculating a median in a spatial window for robustness. For phenomena where the time of repetition varies, the method provides the time of each period which can be further analysed.
@@ 2 Related work
The next group of methods utilise event cameras with a temporal resolution two orders of magnitude higher than RGB cameras. The EB-ASM method [1] measures rotation speeds up to $8500$ RPM by calculating the elapsed time between spikes of distinct polarity events caused by rotating patterns in a manually selected region. However, the method struggles when noise is present, and targets lack high-contrast features. In EV-Tach [19] method, only rotating objects of centrosymmetric shapes (e.g., propeller blades) are assumed, and the method allows for rotation speed measurements of up to $6000$ RPM with a relative mean absolute error as low as $0.3$‰. The Frequency-cam [13] method uses a second-order digital infinite impulse response (IIR) filter to reconstruct per-pixel brightness levels. The time deltas between zero-level crossing are then used to estimate the rotation speed.
In [7], a method for estimation of a table tennis ball spin was proposed. The method uses ordinal time surfaces for tracking, the spin of the ball is estimated from optical flow of events generated by a logo on the ball. The method achieved a spin magnitude mean error of $10.7\pm 17.3$ RPS and a spin axis mean error of $32.9\pm 38.2\degree$ in real time for a flying ball. In [10], accurate frequency estimation of periodic phenomena lacking high-contrast features was demonstrated using a 2D correlation of spatial patterns. However, their method requires manual selection of correlation areas, whereas in our proposed method, no user intervention is necessary.
@@
Avg relative error $37.78$ $\mathbf{19.34}$ $42.96$ $487.98$
Max relative error $\mathbf{3225.00}$ $3362.50$ $7712.50$ $22737.50$
@@ A Simple baseline
    FFT computation – for every pixel and selected polarity, we compute a one-dimensional n-point Discrete Fourier Transform with the Fast Fourier Transform (FFT)  [3] with the event creation time represented by a Dirac pulse.
@@ 4 EEPPR – The proposed method
Our method (see Fig. 1) detects periodic similarities along the time axis of event data in four steps:
@@
    Correlation template selection – the array $\mathbf{E}$ is split into a grid of $K$ spatially non-overlapping areas $\mathbf{A}_{i}$, $i\in\{1,\dots K\}$, of size $W\times W$. In each area, the depth (time axis size) of the correlation template of the area $\mathbf{T}_{i}$ is found by thresholding on a predefined number of events $N$.
@@
    Correlation in 3D – a correlation in 3D of the template $\mathbf{T}_{i}$ is computed with the area $\mathbf{A}_{i}$ (see Fig. 2). The time difference between consecutive peaks in correlation responses is the estimated period of the phenomenon.
@@
Figure 2: Normalised correlation responses of a selected 3D template with $1000$ ms of spatio-temporal event stream. Periodic peaks are highly distinctive, highly regular and indicate a periodic phenomenon.
@@ 5 Experiments
We evaluated all different configurations of parameters (window sizes $W$ of $30$, $45$, $60$ and $75$ and counts of template events $N$ of $100$, $200$, …, $3000$) for the proposed method (see Sec. 4). In all our experiments, the length of the quantisation window $t_{\text{quant}}$ is set to $100$ microseconds. When using window size $W$ set to $45\times 45$ px the method achieves the lowest MRE by a substantial margin and a low maximal relative error compared to results when using other window sizes (see Tab. 2). We noticed a significant improvement in accuracy when using counts of template events $N$ higher than $1500$ (see Tab. 2). Using 3-fold cross-validation on the introduced dataset, the best accuracy is achieved using the window size $W$ set to $45\times 45$ px and the template event count $N$ set to $1800$ events. These parameters are proposed as default parameters of the method.
@@ 5.1 Dataset
Figure 3: Ground-truth rates of experiments in order of appearance in this work.
@@
Event data for three experiments in this work originate from a public dataset [^5] Available at https://docs.prophesee.ai/stable/datasets.html. by Prophesee. For these experiments, we used the EE3P method [10] with manual verification to estimate the GT rate. All other sequences were captured as a part of this work and are publicly available. The final dataset contains sequences of period phenomena with rates ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM), see Fig. 3. Due to the non-constant rates of some periodic phenomena, we restricted our analysis to one-second segments of the event stream. This selection was based on the assumption that the change in rate within such a time frame would be negligible and not significantly impact the accuracy of our measurements.
@@ 5.2 Measuring rotation speed
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Target close-up (d) Aggregated events
@@ 5.3 Measuring frequency of periodic light flashes
The sequence used for this experiment captured the flicker of a mobile phone screen that displayed a white colour at maximum brightness. A top-to-bottom pattern was visible when inspecting the event data, likely due to Pulse Width Modulation (PWM) for backlight control. The GT rate was estimated manually, using the EE3P method and confirmed by findings in a study on the display of this phone [^7] Available at https://bit.ly/notebookcheck-SamsungGalaxyS21UltraReview.. Although the Simple baseline method failed, all other methods achieved a relative error lower than $0.8\%$. The limitation of the proposed method is that it requires the aggregation of events to windows of length $t_{\text{quant}}$ (see Sec. 4). This restricts both the maximal measurable frequency and the theoretical limits of precision of the method. To illustrate, in this experiment, the ground truth period is $\frac{10^{6}}{240}=4166,\bar{6}$ microseconds. Therefore, the closest estimate of the period ($4200$ microseconds) results in an estimated rate of $238.1$ Hz and a relative error of $0.792\%$. The Prophesee method and the FFT baseline method achieved the lowest relative errors ($<0.001$%, see Tab. 3).
@@ 5.4 Measuring vibration frequency of a speaker and a motor
In this experiment, we captured and analysed a sequence capturing a vibrating diaphragm of a speaker with two large low-frequency drivers. An Android application [^8] Application available at https://bit.ly/Frequency-Sound-Generator-APK. precisely controlled the emitted sound (musical note $G_{2}$, $98$ Hz). The event camera captured the vibrating diaphragm of a speaker at approximately 30° angle. The most prominent features in the event stream originate from the vibrating edges of the diaphragm and the manufacturer’s logo located in its centre. The proposed method was the only method that achieved a relative error lower than $0.1\%$. The FFT baseline and Prophesee methods achieved relative error under $0.2\%$, followed by EB-ASM ($1.02\%$) and Frequency-cam ($2.8\%$). The Simple baseline method deviated the most from the reference frequency ($78.571\%$ relative error); see Tab. 3.
@@ 5.6 Discussion
We analysed twelve sequences capturing various periodic phenomena using six different methods. The experiments where the target lacked high-contrast features were shown to be the most challenging. The worst performing methods were the Simple baseline method and the EB-ASM method [1] achieving MRE of $4175.511\%$ and $6848.011\%$ respectively (see Tab. 3). The third lowest mean relative error across all experiments ($32.532\%$) was achieved by the Vibration Estimation method by Prophesee. The Frequency-cam method achieved the second lowest mean relative error (MRE) of $14.083\%$, which is significantly higher than the MRE achieved by the proposed method ($0.1\%$). Note that the Frequency-cam method could not analyse three of the twelve sequences we used in this work, which originated from the public dataset by Prophesee due to an older file encoding type. All other methods were able to analyse all sequences.
@@ 6 Conclusion
A novel contactless method for measuring the rate of periodic phenomena has been presented. The method makes a single assumption that the observed object periodically returns to a known state or position, producing a similar set of events in the event stream. The method, EEPPR, was evaluated on a new dataset of twelve sequences that represent a wide range of periodic phenomena (light flickering, object vibration, rotational speed, and periodic movement), with targets lacking prominent features, being behind transparent materials, and having centrosymmetric shapes, achieving the mean relative error of $0.1\%$. The method works in non-frontal camera placement, measuring frequencies as high as $2000$ Hz (equivalent to 120 000120000120\,000120 000 RPM) even when a significant amount of noise is present.
@@
The method assumes the location where the period phenomenon is occurring is static. Handling of tracking and the template update is future work.
Appendix 0.A Experimental Setup
For the ’Felt disc with a high-contrast line’ experiment (see Sec. 5.2), we captured a power drill spinning a white felt disc marked with a black line, see Fig. 0.A.5.
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.5: Felt disc with a high-contrast line (see Sec. 5.2): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($655\times 655$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
For the ’Fronto-parallel velcro disc’ experiment (see Sec. 5.2), a drill spun the disc at $1266$ RPM with the event camera facing it directly, while a laser tachometer measured speed from the side of the disc where a reflective sticker was attached (Fig. 0.A.6a).
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) RoI
Figure 0.A.6: Fronto-parallel velcro disc (see Sec. 5.2): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($713\times 713$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
In the next experiment (’Velcro disc (side)’, see Sec. 5.2), the velcro disc was captured by the event camera through a glass sheet at $45\degree$ camera angle (see Fig. 4a).
For the following experiment (’High contrast dot’), a sequence from Prophesee’s dataset was used, capturing an orbiting dark dot (see Sec. 5.2). The reference photo was generated using Prophesee’s Event to Video method (see Fig. 0.A.7a).
[Uncaptioned image] [Uncaptioned image]
(a) Target visualisation (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.7: High-contrast dot (see Sec. 5.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($263\times 263$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
This experiment (see Sec. 5.2) used event data from Prophesee’s dataset, capturing a rotating three-blade hand fidget spinner on a flat surface, see Fig. 0.A.8.
[Uncaptioned image] [Uncaptioned image]
(a) Target visualisation (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.8: Hand fidget spinner (see Sec. 5.2): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($239\times 239$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for the next experiment captured the whirling behaviour for evading predators used by the Pholcus phalangioides spider (see Sec. 5.2). A second rotating pattern is visible, caused by the shadow of the spider (see Fig. 0.A.9d,e).
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(a) Physical setup                 (b) Target close-up (c) Region of Interest
[Uncaptioned image] [Uncaptioned image]
(d) Aggregated events (e) XYTime visualisation
Figure 0.A.9: Whirling Pholcus phalangioides (see Sec. 5.2): (a) physical setup; (b) a close-up photo of the target; (d) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($221\times 221$px) shown in (c); (e) events from a 250-millisecond window visualised in spatio-temporal space. Positive events are represented by white colour, and negative events are bright blue.
For the LED experiment (see Sec. 5.3), a circuit with a light-emitting diode (see Fig. 0.A.10a) was used.
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.10: Flashing LED (see Sec. 5.3): (a) physical setup and target detail; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($126\times 126$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for the ’Refreshing mobile phone screen’ experiment (see Sec. 5.3) captured the flicker of a mobile phone screen displaying a white colour at maximum brightness. A moving top-to-bottom pattern was visible when inspecting the event data (see Fig. 0.A.11b).
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.11: Refreshing mobile phone screen (see Sec. 5.3): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within hardware-level region of interest and a time window equal to $\frac{8}{10}$ of the period of the observed phenomenon, highlighted region of interest ($54\times 54$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
In the following experiment (see Sec. 5.4), we captured a vibrating diaphragm of a speaker with two large low-frequency drivers, see Fig. 0.A.12a.
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.12: Speaker diaphragm (see Sec. 5.4): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($631\times 631$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
For the ’Vibrating motor’ experiment (see Sec. 5.4) we used a sequence from Prophesee’s dataset capturing a motor vibrating vertically at $40$ Hz; see Fig. 0.A.13.
[Uncaptioned image] [Uncaptioned image]
(a) Target visualisation (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.13: Vibrating motor (see Sec. 5.4): (a) target visualisation; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($640\times 480$px) within a time window of length equal to the period of the observed phenomenon, highlighted region of interest ($148\times 148$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence used for this experiment (see Sec. 5.5) captured a bicycle chain from a close-up side view (see Fig. 0.A.14a).
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.14: Bike chain from side view (see Sec. 5.5): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured within a hardware-level region of interest and a time window equal to the period of the observed phenomenon, highlighted region of interest ($204\times 204$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.
The sequence for the final experiment (see Sec. 5.5) captured the same bike chain from a top view perspective (see Fig. 0.A.15a).
[Uncaptioned image] [Uncaptioned image]
(a) Physical setup (b) XYTime visualisation
[Uncaptioned image] [Uncaptioned image]
(c) Aggregated events (d) Region of Interest
Figure 0.A.15: Bike chain from a top view (see Sec. 5.5): (a) physical setup; (b) events from a 250-millisecond window visualised in spatio-temporal space; (c) aggregated events captured by the event camera ($1280\times 720$px) within a time window of lenght equal to the period of the observed phenomenon, highlighted region of interest ($180\times 180$px) shown in (d). Positive events are represented by white colour, and negative events are bright blue.

### v3

@@ Abstract
We present a novel method for measuring the rate of periodic phenomena (e.g., rotation, flicker, and vibration), by an event camera, a device asynchronously reporting brightness changes at independently operating pixels with high temporal resolution. The approach assumes that for a periodic phenomenon, a highly similar set of events is generated within a spatio-temporal window at a time difference corresponding to its period. The sets of similar events are detected by a correlation in the spatio-temporal event stream space. The proposed method, EEPPR, is evaluated on a dataset of $12$ sequences of periodic phenomena, i.e. flashing light and vibration, and periodic motion, e.g., rotation, ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM). EEPPR significantly outperforms published methods on this dataset, achieving a mean relative error of $0.1\%$ setting new state of the art. The dataset and codes are publicly available on GitHub.
@@ 1 Introduction
Traditional methods for measuring periodic phenomena often involve contact-based devices such as rotary encoders and mechanical tachometers. These approaches, while established, have inherent limitations: (i) physical contact with the target object can influence its movement and introduce measurement inaccuracies, (ii) contact methods may not be feasible in scenarios with delicate objects, confined spaces, or situations where interfering with the target significantly affects its movement. Laser tachometers offer a less invasive and highly accurate [^1] See https://meters.uni-trend.com/product/ut370-series/#Specifications. alternative but require reflective material (e.g., a sticker) on the target and precise aiming of the laser, as missing the reflective material pass-through results in an inaccurate measurement. Although contactless event-based methods exist, they are either handcrafted for a specific type of periodic phenomena [6] or require user supervision [7, 8].
In this paper, we propose a non-contact method for measuring the properties of any periodic phenomena using an event camera. The input of our method is the event stream, and the output is a single scalar, the rate of the periodic phenomena calculated as a robust point estimate within a spatial window. The method operates remotely, without requiring any modification of the observed object (e.g., a marker). The approach assumes that similar sets of pixel activations of an event camera (events) will occur at regular time intervals in any sequence capturing a periodic phenomenon. These intervals correspond to the period. For phenomena where the time of repetition varies, the method provides the time of each period which can be further analysed.
The method was evaluated on a new dataset of twelve sequences of four types of periodic phenomena: (i) flickering light, (ii) object vibration, (iii) rotation, and (iv) repetitive translational movement. A mean relative error of $0.1\%$ was achieved, which is 140$\approx$ 140$\times$ 140 × lower than the mean relative error of the best available published method (see Tab. 3).
(a) EEPPR method overview
(b) Normalised correlation responses
(c) Ground-truth rates
Figure 1: (a) EEPPR: (i) data captured from an event camera is aggregated into a 3D array, (ii) the array is split into same-sized areas, and in each area a template depth is automatically selected, (iii) a correlation of the template with the event stream is computed in 3D, (iv) a frequency is calculated from the median of time deltas measured between correlation peaks for each window, (v) the output frequency is computed as a median of measurements from all windows. (b) Normalised correlation responses of a selected 3D template with $1000$ ms of spatio-temporal event stream. Periodic peaks are highly distinctive, highly regular and indicate a periodic phenomenon. (c) Ground-truth rates of experiments in order of appearance in this work.
@@ 2 Related work
The next group of methods utilise event cameras with a temporal resolution two orders of magnitude higher than RGB cameras. The EB-ASM method [6] measures rotation speeds up to $8500$ RPM by calculating the elapsed time between spikes of distinct polarity events caused by rotating patterns in a manually selected region. However, the method struggles when noise is present, and targets lack high-contrast features. In EV-Tach [14] method, only rotating objects of centrosymmetric shapes (e.g., propeller blades) are assumed, and the method allows for rotation speed measurements of up to $6000$ RPM with a relative mean absolute error as low as $0.3$‰. The Frequency-cam [7] method uses a second-order digital infinite impulse response (IIR) filter to reconstruct per-pixel brightness levels. The time deltas between zero-level crossing are then used to estimate the rotation speed. In [15], two vibration estimation methods using high-contrast markers are proposed. Both apply the Fast Fourier Transform (FFT) to identify the dominant frequency. The first analyses marker displacement, achieving a maximum relative error of $1.43\%$ for vibrations up to $110$ Hz, the second analyses event generation rate variation within a bounding box, which moves with the marker’s centroid, with a maximum relative error of $1\%$ for vibrations up to $190$ Hz. In [4], a method for estimation of a table tennis ball spin was proposed. The method uses ordinal time surfaces for tracking, the spin of the ball is estimated from optical flow of events generated by a logo on the ball. The method achieved a spin magnitude mean error of $10.7\pm 17.3$ RPS and a spin axis mean error of $32.9\pm 38.2\degree$ in real time for a flying ball. In [16], accurate frequency estimation of periodic phenomena lacking high-contrast features was demonstrated using a 2D correlation of spatial patterns. However, their method requires manual selection of correlation areas, whereas in our proposed method, no user intervention is necessary.
@@
[…]
@@ A Simple baseline
    FFT computation – for every pixel and selected polarity, we compute a one-dimensional n-point Discrete Fourier Transform [17] with the event creation time represented by a Dirac pulse.
@@ 4 EEPPR – The proposed method
Our method (see Fig. 1(a)) detects periodic similarities along the time axis of event data in four steps:
[Uncaptioned image] [Uncaptioned image] [Uncaptioned image] [Uncaptioned image]
(a) Physical setup           (b) XYTime visualisation (c) Target close-up          (d) Aggregated events
@@
    Correlation template selection – the array $\mathbf{E}$ is split into a grid of $K$ spatially non-overlapping areas $\mathbf{A}_{i}$, $i~{}\in~{}\{1,\dots K\}$, of size $W\times W$. In each area, the depth (time axis size) of the correlation template of the area $\mathbf{T}_{i}$ is found by thresholding on a predefined number of events $N$.
@@
    Correlation in 3D – a correlation in 3D of the template $\mathbf{T}_{i}$ is computed with the area $\mathbf{A}_{i}$ (see Fig. 1(b)). The time difference between consecutive peaks in correlation responses is the estimated period of the phenomenon.
@@ 5 Experiments
We evaluated all different configurations of parameters (window sizes $W$ of $30$, $45$, $60$ and $75$ and counts of template events $N$ of $100$, $200$, …, $3000$) for the proposed method (see Sec. 4). In all our experiments, the length of the quantisation window $t_{\text{quant}}$ is set to $100$ microseconds. When using window size $W$ set to $45\times 45$ px the method achieves the lowest MRE by a substantial margin and a low maximal relative error compared to results when using other window sizes (see Tab. 2). We noticed a significant improvement in accuracy when using counts of template events $N$ higher than $1500$ (see Tab. 2). Using 3-fold cross-validation on the introduced dataset, the best accuracy is achieved using the window size $W$ set to $45\times 45$ px and the template event count $N$ set to $1800$ events. We propose these parameter as default.
@@ 5.1 Dataset
Event data for three experiments in this work originate from a public dataset [^5] Available at https://docs.prophesee.ai/stable/datasets.html. by Prophesee. For these experiments, we used the EE3P method [16] with manual verification to estimate the GT rate. All other sequences were captured as a part of this work and are publicly available. The final dataset contains sequences of period phenomena with rates ranging from $3.2$ Hz to 2 kHz (equivalent to 192192192192120 000120000120\,000120 000 RPM), see Fig. 1(c). Due to the non-constant rates of some periodic phenomena, we restricted our analysis to one-second segments of the event stream. This selection was based on the assumption that the change in rate within such a time frame would be negligible and not significantly impact the accuracy of our measurements.
@@ 5.3 Measuring frequency of periodic light flashes
For this experiment we captured the flicker of a mobile phone screen displaying a white colour at maximum brightness. A top-to-bottom pattern is visible in the event data, likely due to Pulse Width Modulation (PWM) for backlight control. The GT rate was estimated manually, using the EE3P method and confirmed by findings in a study on the display of this phone [^7] Available at https://bit.ly/notebookcheck-SamsungGalaxyS21UltraReview.. Although the Simple baseline method failed, all other methods achieved a relative error lower than $0.8\%$. The limitation of the proposed method is that it requires the aggregation of events to windows of length $t_{\text{quant}}$ (see Sec. 4). This restricts both the maximal measurable frequency and the theoretical limits of precision of the method. To illustrate, in this experiment, the ground truth period is $\frac{10^{6}}{240}=4166,\bar{6}$ microseconds. Therefore, the closest estimate of the period ($4200$ microseconds) results in an estimated rate of $238.1$ Hz and a relative error of $0.792\%$. The Prophesee method and the FFT baseline method achieved the lowest relative errors ($<0.001$%, see Tab. 3).
@@ 5.4 Measuring vibration frequency of a speaker and a motor
In this experiment, we captured and analysed vibrating diaphragm of a speaker with two large low-frequency drivers. An Android application [^8] Application available at https://bit.ly/Frequency-Sound-Generator-APK. precisely controlled the emitted sound (musical note $G_{2}$, $98$ Hz). The event camera captured the vibrating diaphragm of a speaker at approximately 30° angle. The most prominent features in the event stream originate from the vibrating edges of the diaphragm and the manufacturer’s logo located in its centre. The proposed method was the only method that achieved a relative error lower than $0.1\%$. The FFT baseline and Prophesee methods achieved relative error under $0.2\%$, followed by EB-ASM ($1.0\%$) and Frequency-cam ($2.8\%$). The Simple baseline method deviated the most from the reference frequency ($78.6\%$ relative error); see Tab. 3.
@@ 5.6 Discussion
We analysed twelve sequences capturing various periodic phenomena using six different methods. The experiments where the target lacked high-contrast features were shown to be the most challenging. The worst performing methods were the Simple baseline method and the EB-ASM method [6] achieving MRE of $4175.5\%$ and $6848.0\%$ respectively (see Tab. 3). The third lowest mean relative error across all experiments ($32.5\%$) was achieved by the Vibration Estimation method by Prophesee. The Frequency-cam method achieved the second lowest mean relative error (MRE) of $14.1\%$, which is significantly higher than the MRE achieved by the proposed method ($0.1\%$). Note that the Frequency-cam method could not analyse three of the twelve sequences we used in this work, which originated from the public dataset by Prophesee due to an older file encoding type. All other methods were able to analyse all sequences.
@@ 6 Conclusion
A novel contactless event-based method, EEPPR, for measuring the rate of various periodic phenomena, such as rotational speed, vibration, flicker and periodic movement, was presented. The method makes only a single assumption that the observed object periodically returns to a known state or position, producing a similar set of events in the event stream. The rate of periodic phenomena is computed by measuring time deltas between peaks detected in correlation responses of an automatically selected template correlated with a spatio-temporal event stream. EEPPR was evaluated on a new proposed dataset of twelve sequences that represent a wide range of periodic phenomena (light flickering, object vibration, rotational speed, and periodic movement), including targets lacking prominent features, being behind transparent materials, or having centrosymmetric shapes, achieving the mean relative error of $0.1\%$ setting new state of the art. The method works in non-frontal camera placement, measuring frequencies as high as $2000$ Hz (equivalent to 120 000120000120\,000120 000 RPM) even when a significant amount of noise is present. The dataset and code are publicly available on GitHub.
@@
The method assumes the periodic phenomenon undergoes only periodic movement without additional motion. Handling of tracking and the template update is future work. Aggregating events within fixed intervals $t_{\text{quant}}$ restricts both the maximal measurable frequency and the theoretical limits of precision. Reducing the aggregation length $t_{\text{quant}}$ mitigates this at a cost of higher computation time. Future work will explore sparse data correlation to avoid aggregation.
Acknowledgements
The research reported in this paper has been funded by BMK, BMAW, and the State of Upper Austria within the SCCH competence center INTEGRATE (grant no. 892418) part of the FFG COMET Competence Excellent Technologies Programme, and by the Czech Technical University in Prague grant No. SGS23/173/OHK3/3T/13.